The shelf life of a particular dairy product is normally distributed with a mean of 12 days and a standard deviation of 3 days. About what percent of the products last between 9 and 12 days?

1
Expert's answer
2022-06-06T03:23:26-0400
"P(9\le X\le 12)=P(Z\le\dfrac{12-12}{3})"

"-P(Z<\dfrac{9-12}{3})=P(Z\le 0)-P(Z<-1)"

"=0.5-0.1586553=0.341345"